Pete, Rob, Alberto,
yesterday we thought that some of the problems we were having in locking the IFO might be related to a change of the length of the mode cleaner. So today we decided to measure it again.
We followed the Sigg-Frolov technique (see 40m Wiki, Waldman, Fricke). For the record, the MC_AO input corresponds to IN2 on the MC Servo board.
We obtained: L = 27.092 +/- 0.001 m
From the new measurement we reset the frequencies of the Marconis to the following values:
33196450 Hz
132785800 Hz
165982250 Hz
199178700 Hz
|